OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2).

OpenStudy (whpalmer4):

First, find the slope of your line \[3x-y=-3\]You could do this by solving for \(y\), which will give you the equation in slope-intercept form: \[y = mx+b\]The slope of the line will be \(m\). Then, as parallel lines have equal slopes, plug that value of \(m\) and the known point \((x_1,y_1) = (4,-2)\) into the point-slope formula: \[y-y_1 = m(x-x_1)\]and rearrange into whatever form you see fit.

OpenStudy (anonymous):

Would I replace the X and Y values given, (4, -2)? Like: \[3(4) - ^{-}2 = ^{-}3?\] @whpalmer4

OpenStudy (whpalmer4):

I can't make any sense of what you just wrote... First step: solve \[3x-y=-3\] for \(y\) on the left side, all alone. What do you get?

OpenStudy (whpalmer4):

You're still going to have numbers and a letter on the right side

OpenStudy (anonymous):

3x - y = -3 + y = + y 3x = y - 3 +3 = + 3 3x + 3 = y y = 3x + 3 I simplified the equation for you now plot that on a graph... now let's find the second equation.

OpenStudy (anonymous):

\[3x = -3 + y?\]

OpenStudy (anonymous):

That was in response to @whpalmer4

OpenStudy (anonymous):

y - (-2) = 3(x - 4) y + 2 = 3x - 12 -2 = -2 y = 3x - 10

OpenStudy (anonymous):

Okay. so now my two equations are \[y = 3x + 3\] and \[y = 3x - 10\]

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

Let me check them first.

OpenStudy (anonymous):

y - (-2) = 3(x - 4) y + 2 = 3x - 12 -2 = -2 y = 3x - 14

OpenStudy (anonymous):

y = 3x + 3 and y = 3x - 14

OpenStudy (anonymous):

Right now I'm practically clueless right now. Here's the answers: \[y = -\frac{ 1 }{ 3 }x - 6\]\[y = -\frac{ 1 }{ 3 }x - 14\]\[y = 3x - 6\] \[y = 3x - 14\]

OpenStudy (anonymous):

Yup, I was right y = 3x - 14.

OpenStudy (anonymous):

This is proof of it being correct if you need proof.

OpenStudy (anonymous):

Okay, thank you!

OpenStudy (anonymous):

Your Welcome. :D

OpenStudy (anonymous):

What do I open that with?

OpenStudy (whpalmer4):

Yes, \[y=3x-14\] is correct. I'll go through the steps I suggested: \[3x-y=-3\]solve for \(y\):\[3x-3x-y=-3-3x\]\[-y=-3-3x\]multiply by -1\[y = 3+3x\]compare with slope-intercept form \(y = mx+b\) and it is clear that \(m = 3\) Now plug known point + known slope into formula: \[y -(-2) = 3(x-4)\]Rearrange to have \(y\) alone on left:\[y = 3(x-4)+(-2)\]\[y = 3x-12-2 = 3x-14\]

OpenStudy (anonymous):

Thank you both so much